IF GNP=500 CAPITAL COSNUMPTION ALLOWANCE=50 INDIRECT TAXES =60 SUBSIDIES=10 CORPORATE PROFITS= 70 DIVIDENDS= 15 GOVERMENT TRANSFER PAYMENTS=25 PERSONAL DIRECT TAXES=70 PERSONAL CONSUMPTION EXPENDITURE= 275 WHAT IS NNP

NATIONAL INCOME , PERSONAL INCOME, PEROSNAL DISPOSABLE INCOME PEROSNAL SAVINGS


1
Expert's answer
2021-04-30T07:23:30-0400
"Solution"

"NNP=GNP-indirect\ taxes+subsidies\\
NNP=500-60=440+10=450"


"Personal \ Income=National \ income -Net \ Transfer \ Payments\\
Personal\ Income=500-165=335"



"Personal \ Disposable\ income=capital\ consumption+personal \ expenditure=50+275=325"


"Personal\ Savings=" "personal \ income -personal direct \ taxes =275-70=195"